OpenStudy (rogue):

What would be the formula for the sequence An, for... 0, 2, 0, 2, ... 1/2, 1/5, 1/10, 1/17, ...

OpenStudy (precal):

0,2,0,2,...... I do not believe has a formula. We tend to write formulas if the sequence is arithmetic or geometric.

OpenStudy (precal):

I could be wrong on that, hopefully someone can give you a formula.

OpenStudy (rogue):

Do you know the equation for the second one though? I'm pretty sure its recursive.

OpenStudy (rogue):

Here are the answers if you still want to know. \[\left\{ 1 + (-1)^n \right\}_{n = 1}^{\infty} = 0, 2,0,2,0,...\]\[\left\{ \frac {1}{n^2 + 1} \right\}_{n = 1}^{\infty} = \frac {1}{2}, \frac {1}{5}, \frac {1}{10},...\]

OpenStudy (precal):

yeah, I believe it. Students do have a hard time coming up with those solutions unless they have a lot of experience with series and sequences.

OpenStudy (rogue):

We were just being given our introduction to infinite series that day. This stuff is pretty easy now =)

OpenStudy (precal):

calculus?

OpenStudy (rogue):

Yeah, BC Calc =)